OpenStudy (anonymous):

For the following system, if you isolated y in the second equation to use the Substitution Method, what expression would you substitute into the first equation? 2x + y = 2 x − y = −6

OpenStudy (anonymous):

I think the question is just asking you to solve the second equation for y

OpenStudy (anonymous):

Can you do that?

OpenStudy (anonymous):

−6 − x −6 + x 6 − x 6 + x

OpenStudy (anonymous):

You're going to have to do the thinking. You need to get y by itself. You can add/subtract/multiply/divide the same thing to each side of the equation in order to achieve this.

OpenStudy (anonymous):

there is 2 ways to do this. It's personal preference. I'll do the first step. -y = -6-x

OpenStudy (anonymous):

They want you to substitute x − y = −6 into the second equation. To do this, you solve for y. x - y = -6 Subtract x from both sides. - y = -x - 6 Divide both sides by -1. y = x + 6

OpenStudy (anonymous):

All in all, you would substitute x + 6 into the first equation. So that would be your answer. (: